# FastVideoEdit: Leveraging Consistency Models for Efficient Text-to-Video Editing

### v1

@@ Abstract
Keywords:
Video Editing Diffusion Models Consistency Models
@@
1 Introduction
Diffusion models [15, 35, 14, 1] have gained significant attention due to their remarkable capabilities in text-to-image [31, 35, 15, 17] and text-to-video generation [14, 34, 3, 12, 4]. Leveraging the capabilities of these models, it becomes feasible to manipulate videos [4] based on textual input, holding great potential for various applications in areas such as film production and content creation.
However, the computational cost associated with sequential sampling in diffusion models presents a significant challenge for efficient inference, especially in video editing scenarios where a set of frames need to be processed. Moreover, the absence of high-quality open-source video diffusion models [10, 28, 10] that can generate consistent editing results within a single test time inference, combined with the constraints on video duration of video diffusion models, has led to the adoption of existing image generation models for achieving accurate video editing [2, 29, 11]. To align the distribution between image and video models and perform accurate video editing, some methods employ a test-time one-shot fine-tuning for inflated image generation model on each input video [38, 33, 23, 37, 26]. However, this process further exacerbates the time-consuming nature of the editing process, which makes it impractical for real-time applications.
@@ 2.1 Video Editing with Diffusion Models
The remarkable success of diffusion-based text-to-image [31, 35, 15, 19, 18] and text-to-video generation models [14, 34, 3, 12, 4] has opened up new possibilities for exciting opportunities in text-based image [13, 17] and video editing [10]. Although editing video directly through video diffusion models [10, 28, 10] show high temporal consistency, the challenges associated with extensive video model training, unstable generation quality, and video duration time limit make using inflated off-the-shelf image generation models a preferable choice for video editing, which inflating 2D model to 3D with an additional temporal channel.
@@ 4 Method
The task of video editing can be described as the following: Given an ordered set of $m$ source video frames $\mathcal{I}_{src}=\{I_{src}^{1},I_{src}^{2},...,I_{src}^{m}\}$ and a source prompt $\mathcal{P}_{src}$ describing the source video, we aim to generate an edited video with temporally consistent frames $\mathcal{I}_{edit}=\{I_{edit}^{1},I_{edit}^{2},...,I_{edit}^{m}\}$ according to a target prompt $\mathcal{P}_{tgt}$.
@@ 4.1 Video Reconstruction with Consistency Model
To our knowledge, FastVideoEdit is the first method in video editing that eliminates the need for the DDIM inversion process while simultaneously performing a complete denoising process on individual video frames. To enable direct editing of the source video without the need for the inversion process, we leverage a consistency model inspired by InfEdit [40]. The key idea to reconstruct source latent is to start with randomly sampled reconstruction noise rather than randomly initialized noisy latents. Following the Multistep Consistency Sampling in Eq 3, we sample a noise $\varepsilon_{t}^{\text{cons}}$ at each timestep $t$ and the noisy latent $z_{t}^{\text{src}}$ becomes directly tractable when $z_{0}^{\text{src}}$ is given in the editing problem. Instead of denoising the randomly initialized noisy latent $z_{T}^{\text{src}}$, the whole trajectory of $\{z_{t}^{\text{src}}\}$ is obtained directly from the sampled noise trajectory $\{\varepsilon_{t}^{\text{cons}}\}$, and in the reverse direction each $\varepsilon_{t}^{\text{cons}}$ can be used to reconstruct $z_{0}^{\text{src}}$ given $z_{t}^{\text{src}}$. The mappings between $z_{t}^{\text{src}}$ and $\{\varepsilon_{t}^{\text{cons}}\}$ given $z_{0}^{\text{src}}$ are given by:
@@
$\varepsilon^{\text{cons}}_{t}$ $=(z_{t}^{\text{src}}-\sqrt{{\alpha}_{t}}z_{0}^{\text{src}})/\sqrt{1-{\alpha}_{t}}.$
@@
Figure 2: Overview of FastVideoEdit. Our model directly denoises three branches of batch frames using three attention control methos: CF-Masa, Re-CA and Bg-Masa. The model uses batch consistency sampling (BCS) with LCMs to improve efficiency, background latent replacement to align editing content with source video and TokenFlow propagation to further improve temporal consistency.
@@ 4.3.1 Cross-Frame Mutual Self-Attention
$\text{attn}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d}}V).$ (10)
@@ 4.3.2 Re-weighted Cross Attention
4.4 Background Preservation via Latent Replacement
There is a trade-off in existing video editing methods between the editing effect of foreground objects and content preservation of background. Changing the attributes of an object in the foreground usually makes the background more consistent with the change. This is because the control methods that are applied to the forward process are not strict control over the latent space. Therefore the change of tokens in the target prompt also influences irrelevant regions of editing latent through attention mechanisms. Compared with state-of-the-art video editing methods, a significant advantage of FastVideoEdit is the accuracy of foreground editing. This is shown in both quantative and qualitative results in Sec. 5. We achieve this by multiple designs of FastVideoEdit. Consistent initial latents and noise in Batch Consistency Sampling algorithm and attention control both provide faithful editing concerning the source video. In addition to this, we propose further background preservation strategies to enhance the faithfulness of the edited content to the source content. Specifically, we propose to simultaneously denoise a background branch that maintains the structure information of the editing branch while aligning content with the source branch. Based on the background branch, we additionally propose a latent replacement algorithm that replaces the background part in the editing latent with the corresponding part in the background latent.
@@ 4.4.1 Background Branch
4.4.2 Latent Replacement
At the end of each denoising step, we employ the latent replacement operation to replace the background region of the editing latent with the corresponding region of the source latent. The region is determined by computing the relative region from a cross-attention map. Specifically, given a cross-attention map $(A^{\text{edit}})_{m\times n}$, we obtain a replacement map $(M^{\text{edit}})_{m}$ where $m$ is the sequence length of the attention map or the size of the feature map, and $n$ is the number of tokens in $\mathcal{P}_{tgt}$. The replacement map is computed as follows:
@@ 4.4.2 Latent Replacement
Intuitively, the replacement map has $1$ at positions where the edited tokens receive high attention scores among all the tokens, and 00 anywhere else. In practice, $A^{\text{edit}}$ is obtained by averaging among all the cross-attention maps of the same size in a fixed resolution level. The replaced edited latent at the end of denoising step $t$ is:
@@ 4.5 Frame Consistency with Tokenflow
$\mathcal{F}_{\gamma}(\mathbf{T}_{base},i,p)=w_{i}\cdot\phi(z^{i+})[\gamma^{i+}[p]]\;+\;(1-w_{i})\cdot\phi(z^{i-})[\gamma^{i-}[p]].$ (17)
In practice, Tokenflow is a plug-and-play operation that can be applied after the self-attention layer. It replaces the original output of spatial features $\phi(z^{i})$ of the original frame latent with the features of weighted sum of two adjacent key frames $\{\mathcal{F}_{\gamma}(\mathbf{T}_{base},i,p)\}_{p}$.
@@
5:     Batched source latents $\mathcal{Z}_{0}^{\text{src}}=\{z_{0}^{\text{src},(i)}\ |\ 1\leq i\leq m\}$
@@
13:     Denoise three branches $\{\varepsilon_{\theta}(\{z^{\text{src}}_{\tau_{n}},z^{\text{edit}}_{\tau_{n}},z^{\text{bg}}_{\tau_{n}}\},{\tau_{n}},\{\mathcal{P}_{src},\mathcal{P}_{tgt},\mathcal{P}_{src})\};\mathbf{T}_{\text{base}}\}$
14:     Update $\mathcal{Z}^{\text{src}}_{\tau_{n+1}}$ using Eq 8
15:     Update $\mathcal{Z}^{\text{edit}}_{0}$ and $\mathcal{Z}^{\text{bg}}_{0}$ using Eq 9
16:     Sample new reconstruction noise $\{\varepsilon_{\tau_{n+1}}^{\text{cons}}\}\sim\mathcal{N}(\bm{0},\bm{I})$
17:     Update $\mathcal{Z}^{\text{edit}}_{\tau_{n+1}}$ and $\mathcal{Z}^{\text{bg}}_{\tau_{n+1}}$ using Eq 8
18:     Replace latents $\mathcal{Z}^{\text{edit}}_{\tau_{n+1}}$ using Eq 14 and 15
19:end for
20:Output: $\mathcal{Z}_{0}^{\textrm{edit}}$
@@ 5.1 Evaluation Benchmark and Metrics
Following previous work [29, 11], we evaluate the temporal consistency of our approach by utilizing clip similarity [30] among frames (‘Tem-Con’). Additionally, we measure the frame-wise editing accuracy through two metrics. ‘Txt-Sim’ for clip similarity between the embeddings of text and image and ‘Clip-Acc’ for the percentage of frames where the edited image has a higher CLIP similarity to the target prompt compared to the source prompt. Furthermore, as an additional evaluation metric, we measure the time consumption of editing $32$ frames’ video using FastVideoEdit and previous methods in both the inversion and forward processes to evaluate the speed.
@@ 5.2 Quantitative Comparison
In Tab. 1 we compare FastVideoEdit with two additional conditional constraints incorporating methods Rerender [41] and Text2Video-Zero [22] as well as three dual-branch methods FateZero [29], Pix2Video [6], and TokenFlow [11].
@@
Table 1: Comparison of FastVideoEdit with previous video editing methods. Bold indicates the best. Underline indicates the second best.
@@
Tem-Con Txt-Sim Clip-Acc Inversion Forward Sum
Rerender [41] 95.7 25.0 48.5 - 174.3 174.3
Text2Video-Zero [22] 96.9 27.1 70.7 - 131.0 131.0
FateZero [29] 95.7 24.9 35.8 233.7 347.0 581.7
Pix2Video [6] 96.0 27.5 68.5 185.3 213.0 399.3
TokenFlow [11] 96.5 25.5 54.7 176.5 115.9 292.4
Ours 96.5 27.7 71.1 - 61.7 61.7
@@ 5.3 Qualitative Comparison
The results show that FastVideoEdit effectively performs video editing aligned with the text prompt while preserving the essential content of the source video. Through attention control, latent replacement, and leveraging the preservation ability of the consistency model, FastVideoEdit successfully performs video foreground editing while preserving the background. This approach enables targeted editing of the foreground elements in the video while ensuring that the background remains intact. By selectively focusing on specific regions of interest and employing latent replacement techniques, FastVideoEdit achieves accurate and consistent editing results, maintaining the integrity of the background content. It is worth noting that FastVideoEdit achieves superior performance compared to other methods while requiring significantly less time. This highlights the efficiency and effectiveness of our approach in delivering high-quality results in a more time-efficient manner.
@@ 5.4 Ablation Study
Tab. 2 shows that without latent replacement the temporal consistency and CLIP accuracy metrics rise, which illustrates that latent replacement protects background but does not help with either temporal consistency or CLIP accuracy. The improvement in background preservation is observed evidently in qualitative results which is not reflected on CLIP metrics. Imposing background preservation prevents the adaption of background to the editing prompt which is negatively reflected on CLIP based similarity evaluation. However, visual observation by eyes can hardly capture the negative impact it causes in terms of content editing. Apart from background preservation designs, the rest of our proposed attention controls achieve better performance in all the three metrics, which shows the effectiveness of our proposed methods.
@@ Conclusion.
In this work, we have introduced FastVideoEdit, a zero-shot video editing approach that addresses the computational challenges faced by previous methods. By leveraging the self-consistency property of Consistency Models (CMs), our method eliminates the need for time-consuming inversion or additional condition extraction steps. We have also introduced a novel approach for maintaining background preservation via latent replacement, which simultaneously denoises a background branch while imposing self-attention control from the source and editing branches. Our experimental results demonstrate the superior performance of FastVideoEdit in terms of editing quality while requiring significantly less time for video editing tasks.
@@
However, FastVideoEdit still has some limitations: (1) FastVideoEdit may require tuning its hyperparameters to achieve optimal performance on each video. This dependency on hyperparameter adjustment adds complexity to the editing process and may require expertise or extensive experimentation to achieve satisfactory results. (2) While FastVideoEdit demonstrates state-of-the-art performance in video editing, there is no guarantee of success for every editing case. The effectiveness of the approach may vary depending on factors such as input data quality, the complexity of the editing task, and the suitability of chosen hyperparameters. (3) The performance of FastVideoEdit relies on the quality and capabilities of the underlying consistency models. In our future work, we are committed to making further improvements to address the challenges highlighted in our research.
Possible negative social impact.
Video editing approaches may pose privacy risks if used to alter videos without appropriate consent or to create content that invades someone’s privacy. Moreover, the convenience and speed offered by FastVideoEdit may inadvertently encourage irresponsible editing practices, leading to ethical dilemmas in areas such as journalism, entertainment, and personal communication. Addressing these possible negative social impacts and continuously improving our models are key focuses for our future releases.

### v2

@@ Abstract
1 Introduction
@@
Diffusion models [15, 34, 14, 1] have gained significant attention due to their remarkable capabilities in text-to-image [30, 34, 15] and text-to-video generation [14, 33, 3, 12, 4]. Leveraging the capabilities of these models, it becomes feasible to manipulate videos [4] based on textual input, holding great potential for various applications in areas such as film production and content creation.
However, the computational cost associated with sequential sampling in diffusion models presents a significant challenge for efficient inference, especially in video editing scenarios where a set of frames need to be processed. Moreover, the absence of high-quality open-source video diffusion models [10, 27] that can generate consistent editing results within a single test time inference, combined with the constraints on video duration of video diffusion models, has led to the adoption of existing image generation models for achieving accurate video editing [2, 28, 11]. To align the distribution between image and video models and perform accurate video editing, some methods employ a test-time one-shot fine-tuning for inflated image generation model on each input video [37, 32, 22, 36, 25]. However, this process further exacerbates the time-consuming nature of the editing process, which makes it impractical for real-time applications.
@@ 2.1 Video Editing with Diffusion Models
The remarkable success of diffusion-based text-to-image [30, 34, 15] and text-to-video generation models [14, 33, 3, 12, 4] has opened up new possibilities for exciting opportunities in text-based image [13, 17] and video editing [10]. Although editing video directly through video diffusion models [10, 27, 10] show high temporal consistency, the challenges associated with extensive video model training, unstable generation quality, and video duration time limit make using inflated off-the-shelf image generation models a preferable choice for video editing, which inflating 2D model to 3D with an additional temporal channel.
@@ 4 Method
Figure 2: Overview of FastVideoEdit. Our model directly denoises three branches of batch frames using three attention control methods: CF-Masa, Re-CA and Bg-Masa. The model uses batch consistency sampling (BCS) with LCMs to improve efficiency, background latent blending to align editing content with source video and TokenFlow propagation to further improve temporal consistency. The right shaded part elaborates detailed operation of using batch consistency sampling to estimate noise in editing branch and background branch.
The task of video editing can be described as the following: Given an ordered set of $m$ source video frames $\mathcal{I}_{\text{src}}=\{I_{1}^{\text{src}},I_{2}^{\text{src}},...,I_{m}^{\text{src}}\}$ and a source prompt $\mathcal{P}_{\text{src}}$ describing the source video, we aim to generate an edited video with temporally consistent frames $\mathcal{I}_{\text{edit}}=\{I_{1}^{\text{edit}},I_{2}^{\text{edit}},...,I_{m}^{\text{edit}}\}$ according to a target prompt $\mathcal{P}_{\text{tgt}}$.
@@ 4.1 Video Reconstruction with Consistency Model
To our knowledge, FastVideoEdit is the first video editing framework that eliminates the need of the DDIM inversion process while simultaneously performing a complete denoising process on individual video frames. To enable direct editing of the source video without the need for the inversion process, we leverage a consistency model inspired by InfEdit [39]. The key idea to reconstruct source latent is to start with randomly sampled reconstruction noise rather than randomly initialized noisy latents. Following the Multistep Consistency Sampling in Eq 3, we sample a noise $\varepsilon_{t}^{\text{cons}}$ at each timestep $t$ and the noisy latent $z_{t}^{\text{src}}$ becomes directly tractable when $z_{0}^{\text{src}}$ is given in the editing problem. Instead of denoising the randomly initialized noisy latent $z_{T}^{\text{src}}$, the whole trajectory of $\{z_{t}^{\text{src}}\}$ is obtained directly from the sampled noise trajectory $\{\varepsilon_{t}^{\text{cons}}\}$, and in the reverse direction each $\varepsilon_{t}^{\text{cons}}$ can be used to reconstruct $z_{0}^{\text{src}}$ given $z_{t}^{\text{src}}$. The mappings between $z_{t}^{\text{src}}$ and $\{\varepsilon_{t}^{\text{cons}}\}$ given $z_{0}^{\text{src}}$ are given by:
@@
$\varepsilon^{\text{cons}}_{t}$ $=(z_{t}^{\text{src}}-\sqrt{{\alpha}_{t}}z_{0}^{\text{src}})/\sqrt{1-{\alpha}_{t}},$
@@ 4.3.1 Cross-Frame Mutual Self-Attention
$\text{attn}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d}})V.$ (10)
@@ 4.3.2 Re-weighted Cross Attention
4.4 Background Preservation via Latent Blending
There is a trade-off in existing video editing methods between the editing effect of foreground objects and content preservation of background. Changing the attributes of an object in the foreground usually makes the background more consistent with the change. This is because the control methods that are applied to the forward process are not strict control over the latent space. Therefore the change of tokens in the target prompt also influences irrelevant regions of editing latent through attention mechanisms. Compared with state-of-the-art video editing methods, a significant advantage of FastVideoEdit is the accuracy of foreground editing. This is shown in both quantative and qualitative results in Sec. 5. We achieve this by multiple designs of FastVideoEdit. Consistent initial latents and noise in Batch Consistency Sampling algorithm and attention control both provide faithful editing concerning the source video. In addition to this, we propose further background preservation strategies to enhance the faithfulness of the edited content to the source content. Specifically, we propose to simultaneously denoise a background branch that maintains the structure information of the editing branch while aligning content with the source branch. Based on the background branch, we additionally propose a latent blending algorithm that replaces the background part in the editing latent with the corresponding part in the background latent.
@@ 4.4.1 Background Branch
4.4.2 Latent Blending
At the end of each denoising step, we employ the latent blending operation to replace the background region of the editing latent with the corresponding region of the source latent. The region is determined by computing the relative region from a cross-attention map. Specifically, given a cross-attention map $(A^{\text{edit}})_{m\times n}$, we obtain a blending map $(M^{\text{edit}})_{m}$ where $m$ is the sequence length of the attention map or the size of the feature map, and $n$ is the number of tokens in $\mathcal{P}_{tgt}$. The blending map is computed as follows:
@@ 4.4.2 Latent Blending
Intuitively, the blending map has $1$ at positions where the edited tokens receive high attention scores among all the tokens, and 00 anywhere else. In practice, $A^{\text{edit}}$ is obtained by averaging among all the cross-attention maps of the same size in a fixed resolution level. The blended edited latent at the end of denoising step $t$ is:
@@ 4.5 Frame Consistency with Tokenflow
$\mathcal{F}_{\gamma}(\mathbf{T}_{base},i,p)$ $=w_{i}\cdot\phi(z^{i+})[\gamma^{i+}[p]]$ (17)
$+(1-w_{i})\cdot\phi(z^{i-})[\gamma^{i-}[p]].$
In general, Tokenflow is a plug-and-play operation that can be applied after the self-attention layer. It replaces the original output of spatial features $\phi(z^{i})$ of the original frame latent with the features of weighted sum of two adjacent key frames $\{\mathcal{F}_{\gamma}(\mathbf{T}_{base},i,p)\}_{p}$.
@@
5:     Source latents $\mathcal{Z}_{0}^{\text{src}}=\{z_{0}^{\text{src},(i)}\ |\ 1\leq i\leq m\}$
@@
13:     Denoise three branches:
14:     $\{\varepsilon_{\theta}(\{z^{\text{src}}_{\tau_{n}},z^{\text{edit}}_{\tau_{n}},z^{\text{bg}}_{\tau_{n}}\},{\tau_{n}},\{\mathcal{P}_{src},\mathcal{P}_{tgt},\mathcal{P}_{src})\};\mathbf{T}_{\text{base}}\}$
15:     Update $\mathcal{Z}^{\text{src}}_{\tau_{n+1}}$ using Eq 8
16:     Update $\mathcal{Z}^{\text{edit}}_{0}$ and $\mathcal{Z}^{\text{bg}}_{0}$ using Eq 9
17:     Sample reconstruction noise $\{\varepsilon_{\tau_{n+1}}^{\text{cons}}\}\sim\mathcal{N}(\bm{0},\bm{I})$
18:     Update $\mathcal{Z}^{\text{edit}}_{\tau_{n+1}}$ and $\mathcal{Z}^{\text{bg}}_{\tau_{n+1}}$ using Eq 8
19:     Replace latents $\mathcal{Z}^{\text{edit}}_{\tau_{n+1}}$ using Eq 14 and 15
20:end for
21:Output: $\mathcal{Z}_{0}^{\textrm{edit}}$
@@ 5.1 Evaluation Benchmark and Metrics
Following previous work [28, 11], we evaluate the temporal consistency of our approach by utilizing clip similarity [29] among frames (‘Tem-Con’). Additionally, we measure the frame-wise editing accuracy through two metrics. ‘Txt-Sim’ for clip similarity between the embeddings of text and image and ‘Clip-Acc’ for the percentage of frames where the edited image has a higher CLIP similarity to the target prompt compared to the source prompt. Consistent with previous research [38], we acknowledge that automated metrics can be noisy and that human evaluation is more reliable. We recognize that automatic metrics may even exhibit a lack of correlation or potentially an inverse correlation with human evaluation results. Therefore, similar to the questions used in Text Guided Video Editing Competition [38], we conduct a user study involving 20 human annotators to evaluate the 76 videos. The annotators rank each group of videos from best to worst according to four aspects: (1) preservation of essential content in the source video (‘P’), (2) video generation quality (‘Q’), (3) temporal consistency among frames (‘C’), and (4) alignment between the text and video (‘T’). Furthermore, as an additional evaluation metric, we measure the time consumption of editing $32$ frames’ video using FastVideoEdit and previous methods in both the inversion and forward processes to evaluate the speed.
@@ 5.2 Quantitative Comparison
In Tab. 1 we compare FastVideoEdit with two additional conditional constraints incorporating methods Rerender [40] and Text2Video-Zero [21] as well as three dual-branch methods FateZero [28], Pix2Video [6], TokenFlow [11], RAVE [19] and DMT [41].
@@
Tem-Con Txt-Sim Clip-Acc P Q C T Inversion Forward Sum
Rerender [40] 95.7 25.0 48.5 5.2 5.8 5.2 4.6 - 174.3 174.3
Text2Video-Zero [21] 96.9 27.1 70.7 7.1 6.7 4.9 4.9 - 131.0 131.0
FateZero [28] 95.7 24.9 35.8 4.5 4.6 2.9 4.8 233.7 347.0 581.7
Pix2Video [6] 96.0 27.5 68.5 4.6 4.8 4.8 4.1 185.3 213.0 399.3
TokenFlow [11] 96.5 25.5 54.7 4.1 3.2 4.3 4.8 176.5 115.9 292.4
RAVE [19] 95.5 26.2 56.8 5.2 4.2 6.5 4.7 69.8 126.4 196.2
DMT [41] 96.2 26.9 70.9 2.4 2.9 3.8 4.2 44.2 363.5 407.5
Ours 96.5 27.7 71.1 2.9 3.8 3.6 3.9 - 61.7 61.7
Table 1: Comparison of FastVideoEdit with previous video editing methods. Bold indicates best. Underline indicates second best.
@@ 5.3 Qualitative Comparison
The results show that FastVideoEdit effectively performs video editing aligned with the text prompt while preserving the essential content of the source video. Through attention control, latent blending, and leveraging the preservation ability of the consistency model, FastVideoEdit successfully performs video foreground editing while preserving the background. This approach enables targeted editing of the foreground elements in the video while ensuring that the background remains intact. By selectively focusing on specific regions of interest and employing latent blending techniques, FastVideoEdit achieves accurate and consistent editing results, maintaining the integrity of the background content. It is worth noting that FastVideoEdit achieves superior performance compared to other methods while requiring significantly less time. This highlights the efficiency and effectiveness of our approach in delivering high-quality results in a more time-efficient manner.
@@ 5.4 Ablation Study
Tab. 2 shows that without latent blending the temporal consistency and CLIP accuracy metrics rise, which illustrates that latent blending protects background but does not help with either temporal consistency or CLIP accuracy. The improvement in background preservation is observed evidently in qualitative results which is not reflected on CLIP metrics. Imposing background preservation prevents the adaption of background to the editing prompt which is negatively reflected on CLIP based similarity evaluation. However, visual observation by eyes can hardly capture the negative impact it causes in terms of content editing. Apart from background preservation designs, the rest of our proposed attention controls achieve better performance in all the three metrics, which shows the effectiveness of our proposed methods.
@@ Conclusion.
In this work, we introduce FastVideoEdit, a zero-shot video editing approach that addresses the computational challenges faced by previous methods. By leveraging the self-consistency property of Consistency Models, our method eliminates the need for time-consuming inversion or additional condition extraction steps. We have also introduced a novel approach for maintaining background preservation via latent blending, which simultaneously denoises a background branch while imposing self-attention control from the source and editing branches. Experimental results demonstrate the superior performance of FastVideoEdit in terms of editing quality while requiring significantly less time for video editing tasks.
@@
FastVideoEdit still has some limitations: (1) FastVideoEdit may require tuning its hyperparameters to achieve optimal performance on each video. This dependency on hyperparameter adjustment adds complexity to the editing process and may require expertise or extensive experimentation to achieve satisfactory results. (2) While FastVideoEdit demonstrates state-of-the-art performance in video editing, there is no guarantee of success for every editing case. The effectiveness of the approach may vary depending on factors such as input data quality and the complexity of the editing task.